OpenStudy (anonymous):

Find an equation of the plane through the point and perpendicular to the given line.

OpenStudy (anonymous):

point: (-2, 7, 10) x = 1 + t, y = 2t, z = 2 - 3t

OpenStudy (anonymous):

From the line you can easily find the plane normal.

OpenStudy (anonymous):

For the point recall we have \(\vec{x}\cdot\vec{N}=\vec{x_0}\cdot\vec{N}\) for any position vector \(\vec{x}\) pointing to a point lying on the plane containing \(\vec{x_0}\) with the normal vector \(\vec{N}\)